Mary Grace invested a certain sum in an account that pays 6% interest and invested P10 000 more in another account that pays 10% interest. If the income from both investments was P34 000, then how much did he invest at each rate?
please give me your table for the solution

Answer by stanbon(75887)

Mary Grace invested a certain sum in an account that pays 6% interest and invested P10 000 more in another account that pays 10% interest. If the income from both investments was P34 000, then how much did he invest at each rate?
--------------------
Interest Equation: 0.06x + 0.10*(x+10,000) = 34000
6x + x + 10000 = 3400000
5x = 3390000
x = 678000 (amt. invested at 6%)
x + 100000 = 688000 (amt. invested at 10%
================
Cheers,
Stan H.